# Detection of evolutionary shifts in variance under an Ornstein–Uhlenbeck model

### v2

@@ Abstract
Abrupt environmental changes can lead to evolutionary shifts in not only the optimal trait value ($\theta$), but also the rate of adaptation ($\alpha$) and the diffusion variance ($\sigma^{2}$) in trait evolution. While several methods exist for detecting shifts in optimal values, few explicitly model shifts in both evolutionary variance and adaptation rates. We use a multi-optima and multi-variance Ornstein-Uhlenbeck (OU) process model (Beaulieu, Jhwueng, Boettiger, and O’Meara, 2012) to describe trait evolution with shifts in both $\theta$ and $\sigma^{2}$ and analyze how covariance between species is affected when shifts in variance occur along the phylogeny. We propose a new method that simultaneously detects shifts in both variance and optimal values by formulating the problem as a variable selection task using an $\ell_{1}$-penalized loss function. Our method is implemented in the R package ShiVa (Detection of evolutionary shifts in variance). Through simulations, we compare ShiVa with methods that only consider shifts in optimal values ($\ell$1ou; PhylogeneticEM). Our method demonstrates improved predictive ability and significantly reduces false positives in detecting optimal value shifts when variance shifts are present. When only shifts in optimal value occur, our method performs comparably to existing approaches. Applying ShiVa to empirical data from cordylid lizards (Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016a), we find that it outperforms $\ell$1ou and PhylogeneticEM, achieving the highest log-likelihood and lowest BIC. evolutionary shift detection, Ornstein-Uhlenbeck model, LASSO, trait evolution, phylogenetic comparative methods
1,2,3Department of Mathematics and Statistics, Dalhousie University, Nova Scotia, Canada
These authors have equal contribution
E-mails: {wn209685, Lam.Ho, tb432381}@dal.ca
@@ 1 Introduction
The process of trait evolution is influenced by environmental conditions. When species experience abrupt environmental changes, their evolutionary processes may also change to adapt to the new environment (Losos, 2011; Mahler, Ingram, Revell, and Losos, 2013a). These changes in the evolutionary process leave signatures in the observed traits of present-day species. By detecting such evolutionary shifts based on observed trait data, we gain valuable insights into historical environmental changes and the evolutionary history of species.
To model the evolutionary process along a phylogenetic tree, researchers commonly use two statistical frameworks: the Brownian Motion (BM) model (Felsenstein, 1985) and the Ornstein-Uhlenbeck (OU) process model (Hansen, 1997). Evolutionary shifts are typically modeled as changes in the parameters of these processes. Much of the existing research focuses on detecting shifts in the current value (under BM) or shifts in the optimal trait value (under OU). Butler and King (2004) formulated the multiple-optima OU model, which allows shifts in the selective optima across different branches of a phylogeny and detects these shifts using hypothesis testing. Uyeda and Harmon (2014) introduced a Bayesian framework to detect such shifts in selective optima. Ho and Ané (2014) highlighted limitations of using AIC and BIC in shift detection tasks, proposing a modified BIC and a forward-backward model selection procedure. Khabbazian et al. (2016) treated the shift detection problem as a variable selection task, applying LASSO to detect shifts in selective optima, implemented in the $\ell$1ou R package. This LASSO-based approach enables efficient shift detection, and they further proposed a new information criterion (pBIC) to account for phylogenetic correlations. Other methods, such as PhylogeneticEM (Bastide, Mariadassou, and Robin, 2017), apply expectation-maximization (EM) to estimate shifts in the optima. Zhang et al. (2024) extended this line of work with the ELPASO package, which uses ensemble variable selection to improve robustness under model violations.
@@
However, these methods primarily focus on detecting shifts in optimal values and do not explicitly detect the shifts in evolutionary diffusion variance. In reality, abrupt environmental changes are likely to influence not only the optimal trait values but also the evolutionary variance, which reflects how much random fluctuation occurs around the evolutionary process. In this paper, the term variance specifically refers to the diffusion variance ($\sigma^{2}$), which quantifies the background rate of stochastic variation in trait evolution. This variance may increase under unstable environments or decrease under stabilizing conditions. Some existing methods allow modeling shifts in evolutionary parameters, including the optimal value ($\theta$), the rate of adaptation ($\alpha$) and the evolutionary variance ($\sigma^{2}$). Pagel and Meade (2006) introduced BayesTraits, which allows shifts in both optimal values and rates of evolution (modeled as $\alpha$). Beaulieu et al. (2012) developed the OUwie package, which supports detecting shifts in both optimal values and the rate of adaptation across branches in the tree. Later, Clavel et al. (2015) introduced mvMorph, which extends evolutionary modeling to multivariate traits, allowing shifts in both the rate of adaptation and evolutionary variance across lineages. More recently, Bastide et al. (2021) developed a flexible Bayesian framework using Hamiltonian Monte Carlo (HMC) sampling, which supports models with shifting evolutionary variances. Finally, Mitov et al. (2020) proposed the $\mathcal{G}_{LInv}$ family of Gaussian phylogenetic models, providing an efficient likelihood-based approach to model variance shifts across branches.
@@
While these tools allow evolutionary shifts to be modeled, they typically require either pre-specified regimes or rely on computationally intensive full Bayesian inference. Furthermore, few methods explicitly focus on detecting both shifts in optimal value and shifts in variance simultaneously using a data-driven variable selection framework. In this paper, we propose a novel LASSO-based method that detects both types of shifts within a unified variable selection framework. Our method uses a multi-optima and multi-variance OU process model, similar to the models used in OUwie (Beaulieu, Jhwueng, Boettiger, and O’Meara, 2012), but innovates by formulating the detection of both shifts in optima and shifts in variance as a single variable selection problem, solvable using LASSO. This approach allows flexible and efficient detection of shifts across the entire tree, without requiring a priori specification of regime boundaries.
We implement this approach in a new R package, ShiVa (Detection of Evolutionary Shifts in Variance). Through simulation studies, we compare ShiVa to existing methods, including $\ell$1ou and PhylogeneticEM, evaluating performance in terms of both shift detection accuracy and predictive performance.
@@
The remainder of the paper is organized as follows. The Trait evolution with shifts in both optimal value and variance section introduces the trait evolution model incorporating shifts in both optimal value and variance, and formally defines the shift detection problem. The Methods section describes the LASSO-based shift detection algorithm. The Simulations section presents simulation results, including comparisons with existing methods. The Case study section illustrates the application of ShiVa to empirical data from cordylid lizards (Broeckhoven, Diedericks, Hui, Makhubo, and Mouton, 2016a). Conclusion concludes with a discussion of the advantages, limitations, and potential extensions of the method.
@@ 2.1 Trait evolution models
The phylogenetic tree is reconstructed from DNA sequences and is assumed to be known in this paper. The phylogenetic tree reveals the correlation structure between trait values of different species. The trait values are correlated based on the shared evolutionary history of species. The trait values of internal nodes are hidden and only trait values of tip nodes can be observed. When $\alpha$ is unknown, a reasonable estimate can be obtained by fitting a null model (a model without shifts) using the phylolm package. Throughout this paper, we focus on estimating the parameters related to shifts, while the trait values of internal nodes remain hidden and only the trait values at the tips of the tree are observed.
Trait evolution models are used to model how the trait values change over time. Brownian Motion and Ornstein–Uhlenbeck are two commonly used models to model the evolution of continuous traits. The BM model is a special case of the OU model. For both models, we let $Y$ denote the vector of observed trait values at the tips, and $Y_{i}$ denote the trait value of taxon $i$. For a single branch, we let $Y_{i}(t)$ denote the trait value at time $t$. These two models assume that conditioning on the trait value of a parent, the evolutionary processes of sister species are independent. We therefore only need to specify the model on one branch. For specifying the model on a single branch, we let $Y(t)$ denote the trait value at time $t$ on a fixed branch.
@@
Brownian motion (BM) was first applied to model the evolution of continuous traits over time by Felsenstein (1985). Under this model, the trait value evolves following a BM process. The process can be written as a stochastic differential equation:
@@
(Hansen, 1997) uses an OU process to model the evolutionary process. A selection force that pulls the trait value toward a selective optimum $\theta$ is included in OU models. An OU process $Y(t)$ is defined by the following stochastic differential equation
@@ 2.2 Evolutionary shifts in optimal value and variance
Butler and King (2004) formulate the multi-optima OU model for adaptive evolution. This model assumes that the optimal value $\theta(t)$ is constant along a branch and may be different between branches. An abrupt change in $\theta(t)$ on a branch is considered an evolutionary shift in optimal value. However, in most previous work, it is assumed that the variance $\sigma^{2}(t)$ is constant throughout the tree. In reality, when a change in the environment happens, not only the optimal values, but also the variances are likely to change. In this paper, we use a multi-optima multi-variance model (Beaulieu, Jhwueng, Boettiger, and O’Meara, 2012) which allows the variances $\sigma(t)$ to change over different branches.
@@
By Itô isometry, the element $\Sigma_{i,j}$ (covariance of $Y_{i}$ and $Y_{j}$) of covariance matrix $\Sigma$ is given by
@@
Where $a$ is the separating point of species $i$ and $j$. For ultrametric trees, the distance between i and j is $2t_{ai}=2t_{aj}=2(T-t_{a})$. For shifts in optimal value, Ho and Ané (2014) showed that the exact location and number of shifts on the same branch are unidentifiable. We therefore also assume that the optimal value is constant along a branch. We let $\theta_{b}$ denote the optimal value on branch $b$, $\Delta\theta_{b}$ denote the change in optimal value on branch $b$ and $\operatorname{path}(\textrm{root},i)$ denote the branches on the path of the phylogenetic tree from the root to the tip $i$. The expectation can be writen as the following equation (Khabbazian, Kriebel, Rohe, and Ané, 2016).
$E(Y_{i})=y_{0}e^{-\alpha T}+(1-e^{-\alpha T})\theta_{0}+\sum_{b\in\operatorname{path}(\textrm{root},i)}(1-e^{-\alpha t_{b}})\Delta\theta_{b}$
Let $\beta_{0}=y_{0}e^{-\alpha T}+(1-e^{-\alpha T})\theta_{0}$ and $\beta_{b}=(1-e^{-\alpha t_{b}})\Delta\theta_{b}$. Then,
$E(Y_{i})=\beta_{0}+\sum_{b}\beta_{b}\mathbf{X_{bi}}$ (1)
@@
where $X_{b}$ is a vector defined by $X_{bi}=0$ if taxon $i$ is not under branch $b$, and $X_{bi}=1$ if the taxon $i$ is under branch $b$.
@@
While shifts in optimal value only affect the expectation of the $Y_{i}$, shifts in variance affect the variance of $Y_{i}$ and the covariance of $Y_{i}$ and $Y_{j}$. In particular, if species $i$ and $j$ separate at time $t_{a}$, and the shifts in variance along the shared evolutionary path occur at times $t_{1},\ldots,t_{m}$, with magnitudes $\Delta\sigma_{1}^{2},\Delta\sigma_{2}^{2},\ldots,\Delta\sigma_{m}^{2}$ respectively, then the covariance $\Sigma_{ij}$ is given by
@@
$\Sigma_{i,j}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left((1-e^{-2\alpha t_{a}})\sigma^{2}_{0}+\sum_{l\in\operatorname{path}(\textrm{root},a)}\left(1-e^{-2\alpha\left(t_{a}-t_{l}\right)}\right)\Delta\sigma^{2}_{l}\right)$
The results of Ho and Ané (2014) about the unidentifiability of shift positions along a branch are not completely applicable for shifts in variance. By examining the covariance between pairs of species under a given branch, we can estimate the covariance at the end of the branch, and from their mean, we can estimate the trait value at the end of the branch. This can result in different likelihoods for different positions of a single shift on a branch, so the position is identifiable. For example, considering the species $i,j,k,l$ on the following tree:
{forest}
for tree= grow = east, parent anchor=east, child anchor=west, align=center, tier/.wrap pgfmath arg=tier #1level(),if n children=0tier=word,font=, [a [a’ [a” [l ][k ]][j ]][i ]]
With a single shift in variance on the branch from a to a’, we have the following equations to solve for the time and magnitude of the shift.
@@
$\Sigma_{jk}-\Sigma_{ij}=\sigma_{0}^{2}/2\alpha(e^{-2\alpha(T-t_{a^{\prime}})}-e^{-2\alpha(T-t_{a})})+\Delta\sigma^{2}/2\alpha(e^{-2\alpha(T-t_{a^{\prime}})}-e^{-2\alpha(T-t_{a})})$
$\Sigma_{kl}-\Sigma_{jk}=\sigma_{0}^{2}/2\alpha(e^{-2\alpha(T-t_{a^{\prime\prime}})}-e^{-2\alpha(T-t_{a^{\prime}})})+\Delta\sigma^{2}/2\alpha(e^{-2\alpha(T-t_{a^{\prime\prime}})}-e^{-2\alpha(T-t_{a^{\prime}})})$
which can easily be seen to be solveable for both $\Delta\sigma^{2}$ and $t$. However, the times, magnitudes and even number of shifts will be unidentifiable if there are two (or more) shifts on the same branch. For example, if there are shifts in variance at times $t_{1}$ and $t_{2}$ on branch $b$, with magnitudes $\Delta\sigma_{1}^{2}$ and $\Delta\sigma_{2}^{2}$ respectively, then for any two species below branch $b$, their covariance will include the term $\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left(1-e^{-2\alpha\left(t_{a}-t_{1}\right)}\right)\Delta\sigma^{2}_{1}+\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left(1-e^{-2\alpha\left(t_{a}-t_{2}\right)}\right)\Delta\sigma^{2}_{2}$. Thus it is not possible to seperate $\Delta\sigma_{1}^{2}$ and $\Delta\sigma_{2}^{2}$, or $t_{1}$ and $t_{2}$. Indeed, in many cases, there is a single shift on branch $b$ that would produce the same covariance matrix.
@@
For model and computational simplicity, we assume that $\sigma^{2}$ is constant over a branch like the assumption for $\theta$. That is, we assume that any shift in variance on a branch occurs at the beginning of that branch. For shifts in mean, this assumption did not limit the space of possible models due to the unidentifiability. However, for shifts in variance, this does restrict the space of models for the trait values at the leaves, and could therefore adversely affect shift detection. We will study whether this adverse effect is cause for concern in our simulation studies.
@@
$\Sigma_{i,j}$ $=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left((1-e^{-2\alpha t_{a}})\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}(1-e^{-2\alpha\left(t_{a}-t_{b}\right)})\Delta\sigma^{2}_{b}\right)$
@@
$\qquad+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha T}}{2\alpha}\left(1-e^{2\alpha t_{b}}\right)\Delta\sigma^{2}_{b}$ (2)
For the ancestral state at the root node, two different assumptions are commonly used, fixed value or stationary distribution. We assume that the ancestral state at the root is a fixed value for the above process. Equation 2.2 is the covariance of node $i$ and $j$ for the OU model with fixed root. For the OU model, the ancestral state at the root is also often assumed to have the stationary distribution. In this case, the covariance of the root node is $\sigma_{0}^{2}/(2\alpha)$ (Ho and Ané, 2013). In this case, the covariance with shifts can be written as:
$\Sigma_{i,j}$ $=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left(1-e^{-2\alpha t_{a}}\right)\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}(1-e^{-2\alpha t_{a}})\Delta\sigma^{2}_{b}$

### v3

@@ Abstract
Sudden changes in environmental conditions can lead to evolutionary shifts not only in the optimal trait value ($\theta$), but also in the diffusion variance ($\sigma^{2}$) under the Ornstein-Uhlenbeck (OU) model. While several methods have been developed to detect shifts in optimal values, few explicitly account for concurrent shifts in both evolutionary variance and diffusion variance. We use a multi-optima and multi-variance OU model to describe trait evolution with shifts in both $\theta$ and $\sigma^{2}$ and analyze how covariance between species is affected when shifts in variance occur along the phylogeny. We propose a new method that simultaneously detects shifts in both variance and optimal values by formulating the problem as a variable selection task using an $\ell_{1}$-penalized loss function. Our method is implemented in the R package ShiVa (Detection of evolutionary shifts in variance). Through simulations, we compare ShiVa with existing methods that can automatically detect evolutionary shifts under the OU model ($\ell$1ou, PhylogeneticEM, and PCMFit). Our method demonstrates improved predictive ability and significantly reduces false positives in detecting optimal value shifts when variance shifts are present. When only shifts in optimal value occur, our method performs comparably to existing approaches. We apply ShiVa to empirical data on floral diameter in Euphorbiaceae and buccal morphology in Centrarchidae sunfishes.
Key words: evolutionary shift detection, Ornstein-Uhlenbeck model, LASSO, trait evolution, phylogenetic comparative methods
@@ 1 Introduction
The process of trait evolution is influenced by environmental conditions. When species experience abrupt environmental changes, their evolutionary processes may also change to adapt to the new environment (Losos,, 2011; Mahler et al., 2013b, ). These changes in the evolutionary process leave signatures in the observed traits of present-day species. By detecting such evolutionary shifts based on observed trait data, we gain valuable insights into historical environmental changes and the evolutionary history of species.
We typically model evolutionary shifts as changes in the value of the parameters of trait evolution models which assume that traits evolve along a phylogenetic tree according to a continuous time Markov process. The two most popular trait evolution models are the Brownian Motion (BM) model (Felsenstein,, 1985) and the Ornstein-Uhlenbeck (OU) model (Hansen,, 1997). This paper will focus on detecting evolutionary shifts under the OU model, which considers both neural drift and natural selection. Although several researchers have studied this topic, much of the existing work focuses only on detecting shifts in the optimal value of the model (Butler and King,, 2004; Uyeda and Harmon,, 2014; Ho and Ané, 2014a, ; Khabbazian et al.,, 2016; Bastide et al.,, 2017; Zhang et al.,, 2024). In reality, abrupt environmental changes are likely to influence not only the optimal value but also the diffusion variance, reflecting how much random fluctuation occurs around the evolutionary process.
@@
In this paper, the term variance specifically refers to the diffusion variance ($\sigma^{2}$) of the OU model, which quantifies the background rate of stochastic variation in trait evolution. This variance may increase in unstable environments or decrease under stabilizing conditions. Some existing methods allow modeling shifts in evolutionary parameters, including the optimal value ($\theta$), the rate of adaptation ($\alpha$), and the evolutionary variance ($\sigma^{2}$). Pagel and Meade, (2006) introduced BayesTraits, a Bayesian framework that models evolutionary shifts across predefined regimes, supporting changes in both optimal trait values and evolutionary rates (modeled through $\alpha$). Beaulieu et al., (2012) developed the OUwie package, which fits a suite of OU-based models that allow for changes in $\theta$, $\alpha$, and $\sigma^{2}$ between user-specified regimes, using maximum likelihood estimation. Clavel et al., (2015) extended these ideas to multivariate traits with mvMorph, fitting multivariate Brownian motion and OU models that accommodate shifts in evolutionary rates and covariances across lineages, also requiring predefined regime mappings. More recently, Gaboriau et al., (2020) developed JIVE, a multi-platform R package for analyzing both intra- and interspecific trait evolution, allowing flexible modeling of shifts in evolutionary parameters, including diffusion variance, under both continuous and discrete regimes. Bastide et al., (2021) proposed a Bayesian framework that models shifts in evolutionary variance without requiring prior regime specification, using Hamiltonian Monte Carlo (HMC) sampling to jointly infer branch-specific parameters and variance shifts. Several methods specifically aim to detect shifts in diffusion variance under the BM model. Eastman et al., (2011) proposed a Bayesian relaxed BM framework, where branch-specific evolutionary rates ($\sigma^{2}$) are drawn from a lognormal distribution. Grundler et al., (2022) introduced a maximum-likelihood dynamic programming approach to efficiently infer branch-specific rates. For the OU model, Mitov et al., (2019) introduced the PCMFit package, which, to our knowledge, is the only available method capable of automatically detecting shifts in diffusion variance ($\sigma^{2}$) without predefined regimes. However, PCMFit assumes that all model parameters shift together across the same set of regimes. In contrast, our method allows independent shifts in the optimal value ($\theta$) and the diffusion variance ($\sigma^{2}$), providing modeling flexibility by permitting shifts to occur at different locations on the tree.
@@
While these tools allow evolutionary shifts to be modeled, they typically require either pre-specified regimes or rely on computationally intensive full Bayesian inference. Furthermore, few methods explicitly focus on detecting both shifts in optimal value and shifts in variance simultaneously using a data-driven variable selection framework. In this paper, we propose a novel LASSO-based method that detects both types of shifts within a unified variable selection framework. Our method uses a multi-optima and multi-variance OU process model, similar to the models used in OUwie (Beaulieu et al.,, 2012), but innovates by formulating the detection of both shifts in optima and shifts in variance as a single variable selection problem. This approach allows flexible and efficient detection of shifts across the entire tree, without requiring a priori specification of regime boundaries. We implement this approach in a new R package, ShiVa (Detection of Evolutionary Shifts in Variance), which detects both shifts in optimal values and variance under the Ornstein–Uhlenbeck model for ultrametric trees and univariate traits. Through simulation studies, we compare ShiVa to existing methods, including $\ell$1ou, PhylogeneticEM and PCMFit evaluating performance in terms of both shift detection accuracy and predictive performance.
@@
The remainder of the paper is organized as follows. The Trait evolution with shifts in both optimal value and variance section introduces the trait evolution model incorporating shifts in both optimal value and variance, and formally defines the shift detection problem. The Methods section describes the LASSO-based shift detection algorithm. The Simulations section presents simulation results, including comparisons with existing methods. The Case study section illustrates the application of ShiVa to two empirical datasets: floral diameter in the parasitic plant family Euphorbiaceae using the flowerTree phylogeny (Davis et al.,, 2007), and buccal morphology in Centrarchidae sunfishes (Revell and Collar,, 2009).
@@ 2.1 Trait evolution models
The phylogenetic tree is reconstructed from DNA sequences and is assumed to be known in this paper. The phylogenetic tree reveals the correlation structure between trait values of different species. The trait values are correlated based on the shared evolutionary history of species. The trait values of internal nodes are hidden and only trait values of tip nodes can be observed. We assume the phylogenetic tree is ultrametric and the selection force $\alpha$ is fixed throughout the tree. When $\alpha$ is unknown, a reasonable estimate can be obtained by fitting a null (no-shift) model using the phylolm package (Ho and Ané, 2014b, ).
Trait evolution models are used to model how the trait values change over time. Brownian Motion and Ornstein–Uhlenbeck are two commonly used models to model the evolution of continuous traits. For both models, we let $Y$ denote the vector of observed trait values at the tips, and $Y_{i}$ denote the trait value of taxon $i$. For a single branch, we let $Y_{i}(t)$ denote the trait value at time $t$. These two models assume that conditioning on the trait value of a parent, the evolutionary processes of sister species are independent. We therefore only need to specify the model on one branch. For specifying the model on a single branch, we let $Y(t)$ denote the trait value at time $t$ on a fixed branch.
@@
Brownian motion (BM) was first applied to model the evolution of continuous traits over time by Felsenstein, (1985). Under this model, the trait value evolves following a BM process. The process can be written as a stochastic differential equation:
@@
Hansen, (1997) uses an OU process to model the evolutionary process. A selection force that pulls the trait value toward a selective optimum $\theta$ is included in OU models. An OU process $Y(t)$ is defined by the following stochastic differential equation
@@ 2.2 Evolutionary shifts in optimal value and variance
Butler and King, (2004) formulate the multi-optima OU model for adaptive evolution. This model assumes that the optimal value $\theta(t)$ is constant along a branch and may be different between branches. An abrupt change in $\theta(t)$ on a branch is considered an evolutionary shift in optimal value. However, in most previous work, it is assumed that the variance $\sigma^{2}(t)$ is constant throughout the tree. In reality, when a change in the environment happens, not only the optimal values, but also the variances are likely to change. In this paper, we use a multi-optima multi-variance model (Beaulieu et al.,, 2012) which allows the variances $\sigma(t)$ to change over different branches. We first derive the formulas assuming a fixed root, and later extend them to account for a random root.
@@
By Itô isometry, the element $\mathbf{\Sigma}_{i,j}$ (covariance of $Y_{i}$ and $Y_{j}$) of covariance matrix $\mathbf{\Sigma}$ is given by
@@
Where $a$ is the separating point of species $i$ and $j$. For ultrametric trees, $t_{a}$ is the time from the root to the separating point. The distance between i and j is $2t_{ai}=2t_{aj}=2(T-t_{a})$. For shifts in optimal value, Ho and Ané, 2014a showed that the exact location and number of shifts on the same branch are unidentifiable. We therefore also assume that the optimal value is constant along a branch. We let $\theta_{b}$ denote the optimal value on branch $b$, $\Delta\theta_{b}$ denote the change in optimal value on branch $b$ and $\operatorname{path}(\textrm{root},i)$ denote the branches on the path of the phylogenetic tree from the root to the tip $i$. The expectation can be writen as the following equation (Khabbazian et al.,, 2016).
$E(Y_{i})=y_{0}e^{-\alpha T}+(1-e^{-\alpha T})\theta_{0}+\sum_{b\in\operatorname{path}(\textrm{root},i)}(1-e^{-\alpha(T-t_{b})})\Delta\theta_{b}$
Let $\beta_{0}=y_{0}e^{-\alpha T}+(1-e^{-\alpha T})\theta_{0}$ and $\beta_{b}=(1-e^{-\alpha(T-t_{b})})\Delta\theta_{b}$. Then,
$E(Y_{i})=\beta_{0}+\sum_{b}\beta_{b}\mathbf{(X_{b})}_{i}$ (1)
@@
where $\mathbf{X_{b}}$ is a vector defined by $\mathbf{(X_{b})}_{i}=0$ if taxon $i$ is not under branch $b$, and $\mathbf{(X_{b})}_{i}=1$ if the taxon $i$ is under branch $b$, and $t_{b}$ denotes the time from the root to the beginning of branch $b$.
@@
While shifts in the optimal value ($\theta$) affect only the expectation of each trait $Y_{i}$, shifts in variance ($\sigma^{2}$) affect both the variance of $Y_{i}$ and the covariance between traits $Y_{i}$ and $Y_{j}$. This makes the location and magnitude of variance shifts potentially more identifiable from the data than shifts in optima. Suppose species $i$ and $j$ diverged at time $t_{a}$, and along their shared ancestry there are variance shifts at times $t_{1},\ldots,t_{m}$ with magnitudes $\Delta\sigma_{1}^{2},\ldots,\Delta\sigma_{m}^{2}$. Then their covariance $\Sigma_{ij}$ is given by:
$\mathbf{\Sigma}_{ij}=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left((1-e^{-2\alpha t_{a}})\sigma_{0}^{2}+\sum_{l=1}^{m}\left(1-e^{-2\alpha(t_{a}-t_{l})}\right)\Delta\sigma_{l}^{2}\right)$
This expression shows that each variance shift contributes to the total covariance based on its position along the evolutionary path. Therefore, the timing of variance shifts affects the likelihood, and different shift placements can result in different covariance structures. This differs from the case of shifts in optima, where the shift location along a branch is not identifiable (Ho and Ané, 2014a, ). In contrast, variance shifts can, in some cases, be localized — particularly when we have multiple descendant species whose pairwise covariances respond differently depending on when the shift occurred.
To illustrate this, consider the tree below, where species $i$, $j$, $k$, and $l$ descend from a common ancestor $a$. Let there be a single variance shift on the branch from $a$ to $a^{\prime}$:
Then we can derive the following two differences in covariances:
@@
$\mathbf{\Sigma}_{jk}-\mathbf{\Sigma}_{ij}$ $=\frac{1}{2\alpha}\left[\sigma_{0}^{2}\left(e^{-2\alpha(T-t_{a^{\prime}})}-e^{-2\alpha(T-t_{a})}\right)+\Delta\sigma^{2}\left(e^{-2\alpha(T-t_{a^{\prime}})}-e^{-2\alpha(T-t)}\right)\right]$
@@
$\mathbf{\Sigma}_{kl}-\mathbf{\Sigma}_{ij}$ $=\frac{1}{2\alpha}\left[\sigma_{0}^{2}\left(e^{-2\alpha(T-t_{a^{\prime\prime}})}-e^{-2\alpha(T-t_{a})}\right)+\Delta\sigma^{2}\left(e^{-2\alpha(T-t_{a^{\prime\prime}})}-e^{-2\alpha(T-t)}\right)\right].$
These two equations are linearly independent and can be solved for both the shift time $t$ and the shift magnitude $\Delta\sigma^{2}$. Thus, in this configuration, the position of a single variance shift along a branch is identifiable.
However, identifiability breaks down when multiple variance shifts occur along the same branch. Suppose two variance shifts occur at times $t_{1}$ and $t_{2}$ on branch $b$, with magnitudes $\Delta\sigma_{1}^{2}$ and $\Delta\sigma_{2}^{2}$. The contribution to the covariance between any two species descending from branch $b$ will be:
$\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left[\left(1-e^{-2\alpha(t_{a}-t_{1})}\right)\Delta\sigma_{1}^{2}+\left(1-e^{-2\alpha(t_{a}-t_{2})}\right)\Delta\sigma_{2}^{2}\right].$
This combined expression depends only on the sum of contributions from both shifts. Therefore, it is not possible to uniquely identify the individual times or magnitudes of these shifts. In fact, a single shift with an appropriately chosen time and magnitude might reproduce the same covariance effect, making the two-shift model unidentifiable.
@@
For model and computational simplicity, we assume that $\sigma^{2}$ is constant over a branch like the assumption for $\theta$. That is, we assume that any shift in variance on a branch occurs at the beginning of that branch. For shifts in mean, this assumption did not limit the space of possible models due to the unidentifiability. However, for shifts in variance, this does restrict the space of models for the trait values at the leaves, and could therefore adversely affect shift detection. We conducted simulation studies to assess the impact of these assumption violations and found that they generally do not affect detection performance, except in the case of two opposing shifts on the same branch—an edge case where the assumptions matter more. Overall, these assumptions are not a major concern. The corresponding simulation results are presented in the Supplementary Materials.
@@
$\mathbf{\Sigma}_{i,j}$ $=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left((1-e^{-2\alpha t_{a}})\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}(1-e^{-2\alpha\left(t_{a}-t_{b}\right)})\Delta\sigma^{2}_{b}\right)$
@@
$\qquad+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha T}}{2\alpha}\left(1-e^{2\alpha t_{b}}\right)\Delta\sigma^{2}_{b}.$ (2)
For the ancestral state at the root node, two different assumptions are commonly used, fixed value or stationary distribution. We assume that the ancestral state at the root is a fixed value for the above process. Equation 2.2 is the covariance of node $i$ and $j$ for the OU model with fixed root. For the OU model, the ancestral state at the root is also often assumed to have the stationary distribution. In this case, the variance of the root node is $\sigma_{0}^{2}/(2\alpha)$ (Ho and Ané,, 2013). In this case, the covariance with shifts can be written as:
$\mathbf{\Sigma}_{i,j}$ $=\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}\left(1-e^{-2\alpha t_{a}}\right)\sigma^{2}_{0}+\sum_{b\in\operatorname{path}(\textrm{root},a)}\frac{e^{-2\alpha(T-t_{a})}}{2\alpha}(1-e^{-2\alpha t_{a}})\Delta\sigma^{2}_{b}$